OpenStudy (anonymous):

Question 1: If f(x) = x2 - x + 2, then find the value for f(-3). Question 2: Evaluate 2xy - m when x = -4, y = 3, and m = 6 Question 3: Solve: -5 + x/3 = 20

OpenStudy (anonymous):

For Question 1, you should replace x with -3. So (-3)2 - (-3) +2 And you solve that

OpenStudy (anonymous):

(-3)2- (-3) +2 = -6 + 3 + 2 = -1

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

Question 2: Evaluate 2xy - m when x = -4, y = 3, and m = 6. To exactly the same thing (replace every variable with its number). So, 2(-4)(3) - 6 is what you need to solve. I won't do this one for you

OpenStudy (anonymous):

Question 3. -5 + x/3 = 20 +5 -5 +x/3 = +5 + 20 (first combine like terms 5 and 20 to get rid of the five on the left side) x/3 = 25 x/3 * 3 = 25 * 3 (multiply both sides by 3 to get rid of the fraction on the left side) x = 75 (this is your answer)

OpenStudy (anonymous):

Your welcome, any more questions? Giving a medal would be appreciated.